Find the area bounded by the ff. curves and lines:

1. the loop of y2 = x 4 (3 - x)

2. 4y = x 2 – 2lnx, y = 0, x = 1, x = 4.

3. y 2 + 4x – 8 = 0, Y = 2x

4. an arch of y =1/3sin(x/3)



1
Expert's answer
2021-12-24T10:24:37-0500

1.


"y^2 = x ^4 (3 - x), x\leq3"

"y=\pm x^2\sqrt{3-x}"

"x^2\sqrt{3-x}=- x^2\sqrt{3-x}"

"x_1=0, x_2=3"

"Area=A=2\displaystyle\int_{0}^{3}x^2\sqrt{3-x}dx"

"\int x^2\sqrt{3-x}dx"

"u=3-x, x=3-u, dx=-du"

"\int x^2\sqrt{3-x}dx=-\int (3-u)^2\sqrt{u}du"

"=-\int(9u^{1/2}-6u^{3/2}+u^{5/2})du"

"=-(9(\dfrac{2}{3})u^{3/2}-6(\dfrac{2}{5})u^{5/2}+\dfrac{2}{7}u^{7/2})+C"

"=-6(3-x)^{3/2}+\dfrac{12}{5}(3-x)^{5/2}-\dfrac{2}{7}(3-x)^{7/2}+C"

"A=2[-6(3-x)^{3/2}+\dfrac{12}{5}(3-x)^{5/2}-\dfrac{2}{7}(3-x)^{7/2}]\begin{matrix}
 3 \\
 0
\end{matrix}"

"=2(0-(-6(3)^{3/2}+\dfrac{12}{5}(3)^{5/2}-\dfrac{2}{7}(3)^{7/2}))"

"=\dfrac{288\sqrt{3}}{35} ({units}^2)"

2.


"4y = x^ 2 – 2\ln x, y = 0, x = 1, x = 4."

"y=\dfrac{1}{4}x^2-\dfrac{1}{2}\ln x"

"Area=A=\displaystyle\int_{1}^{4}(\dfrac{1}{4}x^2-\dfrac{1}{2}\ln x)dx"

"\int\ln xdx=x\ln x-\int x(\dfrac{1}{x})dx=x\ln x-x+C"

"A=[\dfrac{1}{12}x^3-\dfrac{1}{2}x\ln x+\dfrac{1}{2}x]\begin{matrix}
 4\\
 1
\end{matrix}"

"=\dfrac{16}{3}-4\ln 2+2-\dfrac{1}{12}+0-\dfrac{1}{2}"

"=\dfrac{27}{4}-4\ln 2\ ({units}^2)"

3.


"y=2x=>x=\dfrac{1}{2}y"

"y^ 2+4x-8=0=>x=-\dfrac{1}{4}y^2+2"

"-\dfrac{1}{4}y^2+2=\dfrac{1}{2}y"

"y^2+2y-8=0"

"y_1=-4, y_2=2"

"Area=A=\displaystyle\int_{-4}^{2}(-\dfrac{1}{4}y^2+2-\dfrac{1}{2}y)dy"

"=[-\dfrac{1}{12}y^3+2y-\dfrac{1}{4}y^2]\begin{matrix}
 2 \\
 -4
\end{matrix}"

"=-\dfrac{8}{12}+2(2)-\dfrac{4}{4}-(-\dfrac{-64}{12}+2(-4)-\dfrac{16}{4})"

"=9({units}^2)"


4.


"y=\dfrac{1}{3}\sin(\dfrac{x}{3})"




"y=0=>\dfrac{1}{3}\sin(\dfrac{x}{3})=0"

"x=3\pi n, n\in \Z"

"x_0=0, x_1=3\pi"

"Area=A=\displaystyle\int_{0}^{3\pi}\dfrac{1}{3}\sin(\dfrac{x}{3})dx"

"=[-\cos(\dfrac{x}{3})]\begin{matrix}
 3\pi \\
 0
\end{matrix}=-(-1-1)=2({units}^2)"